# Hydrodynamics of a disk in a thin film of weakly nematic fluid subject to linear friction

### v1

@@ V Finite-element simulations
To validate the analytical results, we carry out numerical simulations using the Gascoigne 3d finite element library [61]. For this purpose, the set of Eqs. (1) and (5) is formulated for $\bm{f}(\bm{r})=\bm{0}$ in variational formulation. Both fields, velocity $\bm{v}(\bm{r})$ and pressure $p(\bm{r})$ are discretized using quadratic finite elements. As this equal-order finite element pair is not satisfying the inf-sup condition, the discretization is stabilized with local projections [62]
$\bm{\nabla}\cdot\bm{v}-\sum_{T\in\Omega_{h}}\delta_{T}\Delta(p-\pi^{(1)}p)=0,$
where $\pi^{(1)}$ is the projection to the space of linear finite elements with the parameter $\delta_{T}=(\alpha+h_{T}^{-2})^{-1}$ that depends on the local grid size $h_{T}$ and on $\alpha$.
@@
We choose $L$ large enough so that the solution decays sufficiently in magnitude when reaching the boundaries. We thus assume $\bm{v}=\bm{0}$ as Dirichlet conditions on the outer boundary. On the disk, $\bm{v}=(V,0)$ or $\bm{v}=(0,V)$ is enforced. For $L=200$, the artificial truncation of the domain is no longer perceptible as long as $\alpha\geq 0.1$. For smaller values of $\alpha$, we must choose $L$ even larger. To keep computing time low, we use local mesh refinement and only refine elements within the square $[-10,10]^{2}$. The local mesh size at the disk is about $h_{T}\approx 0.015$. The resulting linear problems are solved with a GMRES solver [63], preconditioned by a parallel multigrid method [64]. Computing times are just a few seconds per configuration.
To evaluate the drag on the disk, we integrate the viscous stress along the perimeter of the disk, taking into account the direction of motion. We use the Babuška-Miller trick [65] to express these line integrals via volume integrals. This formulation gives the drag with highest accuracy and fourth-order convergence in the mesh size $h_{T}$, see Refs. 66 and 67, (Remark 8.17) for further details.
@@ VI Results and discussion
We thus proceed to assess the adequacy of the perturbative expansion in the analytical theory through rigorous comparison with FEM simulations. To focus on the deviation from isotropic Brinkman effects in scenarios of weak anisotropy, we showcase the image flow field $\bm{v}^{*}$ rather than the complete flow field $\bm{v}$. In Figs. 3, we present streamlines and quiver plots illustrating $\bm{v}^{*}$, for motion parallel to the director, that is, into the $x$-direction. Results are presented for three cases, namely for $c=0.1$ (left column), $c=1$ (middle column), and $c=10$ (right column). Additionally, the parameters $A$ and $\beta$ are varied as $(A,\beta)=(0.5,0)$ (top row) and $(A,\beta)=(0,0.5)$ (bottom row). Again, we present the analytical solutions in the top halves of the plots ($y>0$), while FEM results are shown for comparison in the bottom halves ($y<0$).
@@
Finally, in Fig. 5 we challenge the theory by examining values of $A$ and $\beta$ that we may assume are beyond the range of validity of the perturbative expansion. We keep $c=\alpha_{\parallel}a$ fixed at the magnitude $c=1$. For $(A,\beta)=(-2,0)$ as depicted in panels (a) and (b), the theory tends to underestimate the magnitude of the velocity field $\bm{v}^{*}$ when compared to FEM results. We remark that for negative values of $A$ and $\beta$, the perturbative solution for the image flow field $\bm{v}^{*}$ at the leading order provides the same flow field, except for an opposite sign. Both, translation of the disk (a) parallel and (b) perpendicular to the axis of uniaxiality are considered, with theoretical solutions at the top and on the right-hand side, while FEM results are included at the bottom and on the left-hand side, respectively. The same applies to panels (c) and (d), respectively, where we illustrate the case of $(A,\beta)=(0,-4)$. There, the theory overestimates the magnitude of the velocity field $\bm{v}^{*}$ when compared to FEM results. Still, in all considered cases, there is convincing qualitative agreement concerning the structure of the flow fields and streamlines. Enhanced quantitative agreement between theory and simulations could potentially be achieved by including higher-order terms in the expansion concerning the anisotropy of the fluid, which may be explored in future research.
@@ VII Conclusions
Overall, due to the specific features of two-dimensional flows, we hope to stimulate experimental confirmation. Introducing a globally ordered uniaxially anisotropic fluid layer may be achieved, for instance, in liquid-crystalline films on supporting substrates, if the substrate is rubbed accordingly to orient the nematic director. Incompressibility of the fluid within the plane can be realized by introducing an additional top coverslide to avoid varying thickness of the liquid-crystalline film. Another extension concerns the situation of thin elastic films and membranes. Our plan is to address membranes of uniaxial nematic elastomers [70] as a next step.
@@
The Maple 23 scripts containing the final expressions for the velocity and pressure fields are accessible through the Zenodo repository at [link].

### v2

@@ V Finite-element simulations
To validate the analytical results, we carry out numerical simulations using the Gascoigne 3d finite element library [93]. For this purpose, the set of Eqs. (1) and (5) is formulated for $\boldsymbol{f}(\boldsymbol{r})=\boldsymbol{0}$ in variational formulation. Both fields, velocity $\boldsymbol{v}(\boldsymbol{r})$ and pressure $p(\boldsymbol{r})$ are discretized using quadratic finite elements. As this equal-order finite element pair is not satisfying the inf-sup condition, the discretization is stabilized with local projections [94]
$\boldsymbol{\nabla}\cdot\boldsymbol{v}-\sum_{T\in\Omega_{h}}\delta_{T}\Delta(p-\pi^{(1)}p)=0,$
where $\pi^{(1)}$ is the projection to the space of linear finite elements with the parameter $\delta_{T}=({\alpha_{\parallel}}+h_{T}^{-2})^{-1}$ that depends on the local grid size $h_{T}$ and on ${\alpha_{\parallel}}$.
@@
We choose $L$ large enough so that the solution decays sufficiently in magnitude when reaching the boundaries. We thus assume $\boldsymbol{v}=\boldsymbol{0}$ as Dirichlet conditions on the outer boundary. On the disk, $\bm{v}=(V,0)$ or $\bm{v}=(0,V)$ is enforced. For $L=200$, the artificial truncation of the domain is no longer perceptible as long as ${\alpha_{\parallel}}\geq 0.1$. For smaller values of ${\alpha_{\parallel}}$, we must choose $L$ even larger. To keep computing time low, we use local mesh refinement and only refine elements within the square $[-10,10]^{2}$. The local mesh size at the disk is about $h_{T}\approx 0.015$. The resulting linear problems are solved with a GMRES solver [95], preconditioned by a parallel multigrid method [96]. Computing times are just a few seconds per configuration.
To evaluate the drag on the disk, we integrate the viscous stress along the perimeter of the disk, taking into account the direction of motion. We use the Babuška-Miller trick [97] to express these line integrals via volume integrals. This formulation gives the drag with highest accuracy and fourth-order convergence in the mesh size $h_{T}$, see Refs. 98 and 99, (Remark 8.17) for further details.
@@ VI Results and discussion
We thus proceed to assess the adequacy of the perturbative expansion in the analytical theory through rigorous comparison with FEM simulations. To focus on the deviation from isotropic Brinkman effects in scenarios of weak anisotropy, we showcase the image flow field $\boldsymbol{v}^{*}$ rather than the complete flow field $\boldsymbol{v}$. In Figs. 4, we present streamlines and quiver plots illustrating $\boldsymbol{v}^{*}$, for motion parallel to the director, that is, into the $x$-direction. Results are presented for three cases, namely for $c=0.1$ (left column), $c=1$ (middle column), and $c=10$ (right column). Additionally, the parameters $A$ and $\beta$ are varied as $(A,\beta)=(0.5,0)$ (top row) and $(A,\beta)=(0,0.5)$ (bottom row). Again, we present the analytical solutions in the top halves of the plots ($y>0$), while FEM results are shown for comparison in the bottom halves ($y<0$).
@@
In Fig. 6 we challenge the theory by examining values of $A$ and $\beta$ that we may assume are beyond the range of validity of the perturbative expansion. We keep $c=\alpha_{\parallel}a$ fixed at the magnitude $c=1$. For $(A,\beta)=(-2,0)$ as depicted in panels (a) and (b), the theory tends to underestimate the magnitude of the velocity field $\boldsymbol{v}^{*}$ when compared to FEM results. We remark that for negative values of $A$ and $\beta$, the perturbative solution for the image flow field $\boldsymbol{v}^{*}$ at the leading order provides the same flow field, except for an opposite sign. Both, translation of the disk (a) parallel and (b) perpendicular to the axis of uniaxiality are considered, with theoretical solutions at the top and on the right-hand side, while FEM results are included at the bottom and on the left-hand side, respectively. The same applies to panels (c) and (d), respectively, where we illustrate the case of $(A,\beta)=(0,-4)$. There, the theory overestimates the magnitude of the velocity field $\boldsymbol{v}^{*}$ when compared to FEM results. Still, in all considered cases, there is convincing qualitative agreement concerning the structure of the flow fields and streamlines. Enhanced quantitative agreement between theory and simulations could potentially be achieved by including higher-order terms in the expansion concerning the anisotropy of the fluid, which may be explored in future research.
When comparing the resulting magnitudes of the flow field upon varying the anisotropy parameters $A$ and $\beta$, we observe that the fluid velocity is significantly larger for nonzero values of $\beta$ than for nonzero values of $A$ of equal magnitude. Apparently, changes in $A$ have a less significant impact across the wide range of values of $c$ considered. This trend is also observed when we address the direction of motion. The magnitude of the velocity is larger for motion perpendicular to the director than parallel to it. These observations align with the previously discussed behavior regarding the contribution of each anisotropy parameter to the correction factors in the hydrodynamic mobilities, as long as $c=\alpha_{\parallel}a$ does not grow too large, see Fig. 2.
Finally, we present in Fig. 7 the total flow field $\boldsymbol{v}$ (not only the image flow field $\boldsymbol{v}^{*}$) as obtained from finite element simulations using the parameters from Fig. 6. When compared to the flow in an isotropic medium in Fig. 3 (b) for $c=1$, we observe that for $(A,\beta)=(-2,0)$, see Fig. 7 (a) and (b), the structure of the flow field is not significantly affected by the anisotropy in viscosity. Conversely, the case $(A,\beta)=(0,-4)$ in Fig. 7 (c) and (d) demonstrates a strong impact of the anisotropy in friction on the flow behavior and streamlines. Since in our investigations the contribution of the image solutions to the total flow field is generally small compared to the isotropic component, we do not observe the vortical structures when concentrating on the overall flow.
@@ VII Conclusions
Overall, due to the specific features of two-dimensional flows, we hope to stimulate experimental confirmation. Introducing a globally ordered uniaxially anisotropic fluid layer may be achieved, for instance, in liquid-crystalline films on supporting substrates, if the substrate is rubbed accordingly to orient the nematic director. Incompressibility of the fluid within the plane can be realized by introducing an additional top coverslide to avoid varying thickness of the liquid-crystalline film. Another extension concerns the situation of thin elastic films and membranes. Our plan is to address membranes of uniaxial nematic elastomers [100, 101, 102] as a next step.
Figure 7: Total flow field $\boldsymbol{v}$ for the same parameters as in Fig. 6, where the image flow field $\boldsymbol{v}^{*}$ was illustrated. The results on display are obtained from finite element simulations.
Appendix A List of parameters
In Tab. 2, we summarize and provide an overview of the key parameters introduced in the main text for quick reference.
  Parameter Definition
$\bar{\nu}$ $2\left(\nu_{1}+\nu_{2}-2\nu_{3}\right)$
$\zeta_{1}$ $\left(2\nu_{1}+\nu_{2}-\nu_{4}+\nu_{5}\right)/\nu_{3}-1$
$\zeta_{2}$ $\left(\nu_{2}+\nu_{4}-\nu_{5}\right)/\nu_{3}-1$
$\alpha_{\parallel}^{2}$ $m_{\parallel}^{2}/\nu_{3}$
$\alpha_{\perp}^{2}$ $m_{\perp}^{2}/\nu_{3}$
$A$ $\bar{\nu}/\nu_{3}$
$B$ $1+A\sin^{2}\phi\cos^{2}\phi$
$H$ $\alpha_{\parallel}^{2}\sin^{2}\phi+\alpha_{\perp}^{2}\cos^{2}\phi+k^{2}B$
$w$ $a/\rho$
$\sigma$   $\left.\rho\left(\alpha_{\parallel}^{2}\sin^{2}\phi+\alpha_{\perp}^{2}\cos^{2}\phi\right)^{1/2}\middle/B^{1/2}\right.$
$\beta$ $1-\alpha_{\perp}/\alpha_{\parallel}$
$r$ $\alpha_{\parallel}\rho$
$c$ $\alpha_{\parallel}a$
$b$ $\left(\nu_{1}-\nu_{4}+\nu_{5}\right)/\nu_{3}$
ϵ±subscriptitalic-$\epsilon_{\pm}$ $\left(A\pm 2b\right)/8$
Table 2: List of the main parameters defined and used in the main text.
Appendix B Analytical expression of $\Phi_{mn}$
In the following, we provide the proof for the analytical result listed in Eq. (36). To begin, we define the integral
$g(\sigma,w)=\int_{0}^{\infty}\frac{x\,J_{2m}(x)J_{2n}({w}x)}{x^{2}+\sigma^{2}}\,\,\mathrm{d}x\,,$ (69)
with $m,n\in\mathbb{N}$, $\sigma\in\mathbb{R}$, and $w\in[0,1]$.
Next, we introduce the Hankel function of the first kind of order $2m$ as [85]
$H_{2m}^{(1)}(z)=J_{2m}(z)+iY_{2m}(z)\,.$ (70)
In this expression, $Y_{2m}$ denotes the Bessel function of the second kind of order $2m$. The Hankel function of the first kind, $H_{2m}^{(1)}$, has a branch cut along the negative real axis. On the upper side of this branch cut, for $x>0$, we have
$H_{2m}^{(1)}(-x)=-J_{2m}(x)+iY_{2m}(x)\,.$ (71)
We recall that, for $x>0$,
$J_{\nu}(xe^{i\pi})$ $=e^{i\pi\nu}J_{\nu}(x)\,,$ (72a)
$Y_{\nu}(xe^{i\pi})$ $=e^{-i\pi\nu}Y_{\nu}(x)+2i\cos(\pi\nu)J_{\nu}(x)\,.$ (72b)
Then, since
$\frac{x\,J_{2n}({w}x)}{x^{2}+\sigma^{2}}$ (73)
is an odd function of $x$, we find
$g(\sigma,w)=\frac{1}{2}\operatorname{Re}\int_{-\infty}^{\infty}\frac{x\,H_{2m}^{(1)}(x)J_{2n}({w}x)}{x^{2}+\sigma^{2}}\,\,\mathrm{d}x\,,$ (74)
where the integration from $\infty$- ∞ to 0 is taken along the upper side of the branch cut.
Now, let us define the function $f(z)$ in the complex plane as
$f(z)=\frac{z\,H_{2m}^{(1)}(z)J_{2n}({w}z)}{z^{2}+\sigma^{2}}\,.$ (75)
Performing a series expansion of $f(z)$ around $z=0$, we find that $\lim_{z\to 0}f(z)=0$ for $m\leq n$. Conversely, for $m>n$, $f(z)$ around $z=0$ can be expressed as
$f(z)=\sum_{k=-1}^{-m}a_{2k+1}z^{2k+1}+\mathcal{O}(z)\,.$ (76)
As there are no even negative power terms in the series expansion provided by Eq. (76), we can express
$g(\sigma,w)=\frac{1}{2}\,\operatorname{Re}\,\mathchoice{{\vbox{\hbox{$\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\scriptstyle-$}}\kern-3.25pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.29166pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-1.875pt}}\!\int_{-\infty}^{\infty}f(x)\,\mathrm{d}x\,,$ (77)
with the dashed integral representing the Cauchy principal value.
Let us integrate the function $f(z)$ along a contour composed of the upper side of the branch cut from $-R$ to $-\epsilon$, a small clockwise-oriented semicircle of radius $\epsilon$ centered at the origin denoted by $C_{\epsilon}$, the positive real axis from $\epsilon$ to $R$, and the counterclockwise path along the upper half of the circle $|z|=R$.
Leveraging that as $z\to\infty$
$J_{\nu}(z)$ $\sim\left(\frac{2}{\pi z}\right)^{\frac{1}{2}}\cos\left(z-\frac{\nu\pi}{2}-\frac{\pi}{4}\right)\,,$ (78a)
$H_{\nu}^{(1)}$ $\sim\left(\frac{2}{\pi z}\right)^{\frac{1}{2}}e^{i\left(z-\frac{\nu\pi}{2}-\frac{\pi}{4}\right)}\,,$ (78b)
it follows that
$|f(z)|\sim\frac{2}{\pi}\frac{1}{w^{\frac{1}{2}}|z|^{2}}\,\cos\left(z-\frac{\pi}{4}\right),$ (79)
as $|z|\to\infty$ in the upper half plane. Consequently, as $R$ approaches infinity, the integral vanishes over the upper half of the circle $|z|=R$.
Then, as $\epsilon$ tends to zero, the residue theorem yields
[…]
Using the connecting formulas [85]
$I_{\nu}(z)$ $=e^{\mp\frac{i\pi\nu}{2}}J_{\nu}\left(ze^{\pm\frac{i\pi}{2}}\right)\,,$ (81a)
$K_{\nu}(z)$ $=\frac{i\pi}{2}\,e^{\frac{i\nu\pi}{2}}H_{\nu}^{(1)}\left(ze^{\frac{i\pi}{2}}\right)\,,$ (81b)
the residue at $z=i\sigma$ is calculated and expressed as
$\operatorname{Res}\left\{f(z),i\sigma\right\}=\frac{(-1)^{m+n}}{2i\pi}\,K_{2m}(\sigma)I_{2n}(w\sigma)\,.$ (82)
Now, let us focus on the contour integral along the small clockwise-oriented semicircle of radius $\epsilon$ centered at the origin. Since $f(z)$ is an odd function, we can evaluate this integral using the residue theorem by considering a complementary semicircle of radius $\epsilon$ in the lower plane. This approach leads to
$\int_{C_{\epsilon}}f(z)\,\mathrm{d}z=-i\pi\operatorname{Res}\left\{f(z),0\right\}\,,$ (83)
The negative sign arises from the clockwise direction of the contour integration. Utilizing partial fraction decomposition, we obtain
$\operatorname{Res}\left\{f(z),0\right\}=\operatorname{Res}\left\{h(z,-i\sigma),0\right\}+\operatorname{Res}\left\{h(z,i\sigma),0\right\}\,,$ (84)
where
$h(z,\alpha)=-\frac{1}{2}\frac{H_{2m}^{(1)}(z)J_{2n}(wz)}{\alpha-z}\,.$ (85)
From complex calculus, it is known that, if a function $\phi(z)$ is analytic at $z=\alpha$ and has a pole at $z=a$, then the residue of $\phi(z)/\left(\alpha-z\right)$ at $z=a$ equals the principal part of the Laurent series expansion of $\phi(z)$ about $z=a$, evaluated at $z=\alpha$. Accordingly,
$\operatorname{Res}\left\{f(z),0\right\}$ $=-\frac{1}{2}\,\mathcal{P}H_{2m}^{(1)}(-i\sigma)J_{2n}(-iw\sigma)$
$\quad-\frac{1}{2}\,\mathcal{P}H_{2m}^{(1)}(i\sigma)J_{2n}(iw\sigma)\,.$ (86)
Using the connecting formula [85]
$H_{\nu}^{(1)}(e^{-\frac{i\pi}{2}}z)=2e^{-\frac{i\nu\pi}{2}}I_{\nu}(z)-\frac{2i}{\pi}e^{\frac{i\nu\pi}{2}}K_{\nu}(z)\,$ (87)
together with Eqs. (81) and using $\mathcal{P}I_{2m}(\sigma)I_{2n}(w\sigma)=0$, we obtain
$\operatorname{Res}\left\{f(z),0\right\}=\frac{2}{\pi}\,(-1)^{m+n}\,\mathcal{P}K_{2m}(\sigma)I_{2n}(w\sigma)\,,$ (88)
which is zero when $n\geq m$.
Finally, we combine the results utilizing Eqs. (80), (82), (83), and (88). In this way, we arrive at the analytical solution presented in Eq. (36).
@@ Appendix B Analytical expression of $\Phi_{mn}$
The Maple 23 scripts containing the final expressions for the velocity and pressure fields are accessible through the Zenodo repository using the DOI 10.5281/zenodo.10817264.